James wishes to obtain 800 mL of a 7% solution of boric acid by mixing 4% and 12% solutions. How much of each should be used?
Found 2 solutions by josgarithmetic, MathTherapy:
Answer by josgarithmetic(39617)

No. Boric acid solution will normally be at most, about 5% in water. Is the temperature near or more than 60 C?
http://sti.srs.gov/fulltext/SRNL-STI-2011-00578.pdf

Answer by MathTherapy(10552)

James wishes to obtain 800 mL of a 7% solution of boric acid by mixing 4% and 12% solutions. How much of each should be used?
Let amount of 4% solution to mix, be F
Then amount of 12% solution to mix is: 800 - F
The following equation is thus formed: .04F + .12(800 - F) = .07(800)
.04F + 96 - .12F = 56
.04F - .12F = 56 - 96
- .08F = - 40